please help me convert it
{81^3/4}=27 into logarithm form
3/4 is the power..

Answer by lwsshak3(11628)

please help me convert it
{81^3/4}=27 into logarithm form
3/4 is the power
**
log base 81 (27)=3/4
in other words,
log of 27, base 81=3/4
Base(81) raised to log of number,(3/4)=number(27)